OpenStudy (anonymous):

What is the solution of this linear system? 3x + 5y = -1 2x − 5y = 16

OpenStudy (jack1):

if 3x + 5y = -1 what does 5y = ?

OpenStudy (anonymous):

3x + 5y = -1 2x-5y=16

OpenStudy (jack1):

using only the first equation: if: \[3x + 5y = -1\] what does 5y = ?

OpenStudy (jack1):

yo, you there bro?

OpenStudy (anonymous):

-1?

OpenStudy (jack1):

no, you would subtract 3x from both sides to get 5y by itself so if 3x + 5y = -1 5y = -1 -3x

OpenStudy (anonymous):

lol i thought you were asking me what the number was

OpenStudy (jack1):

ahh, ok now sub that value for 5y into your 2nd equation: so 2x - 5y = 16 ( and 5y = -1 -3x... therefore) 2x - (-1 -3x) = 16

OpenStudy (jack1):

can you solve that equation by adding and subtracting the values to get x = something?

OpenStudy (anonymous):

i need to get the x vaule which is -3x and y vaule

OpenStudy (anonymous):

which is either 2 or -2

OpenStudy (jack1):

yep, and from this equation : 2x - (-1 -3x) = 16 you can get the value for x and if you have the value for x, you can sub it into either equation and it will give you the value for y

OpenStudy (jack1):

give it a shot 2x - (-1 -3x) = 16 2x --1 --3x = 16 2x +1 +3x = 16 5x + 1 = 16 so what does x equal?

OpenStudy (anonymous):

wait i thought x was -3

OpenStudy (jack1):

how did u get x = -3 man?

OpenStudy (anonymous):

is it 2?

OpenStudy (jack1):

you cant just guess, its maths, you have to solve it so: 5x + 1 = 16 5x = 15 so what is x equal to?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

okay i found y =-2

OpenStudy (jack1):

booyah, so if you put x = 3 into the first equation you had: 3x + 5y = -1 so 3 times 3 + 5y = -1 9 + 5y = -1 so what does y equal?

OpenStudy (jack1):

sweet

OpenStudy (anonymous):

thanks bro

OpenStudy (jack1):

props

OpenStudy (jack1):

all good